3 g of Calcium Chloride (CaCl2) is added to 500 mL of water. It all dissolves. Determine the concentration of Chloride ions in the solution. (Hint: there is a trick here. Note: The chemical formula below that shows Calcium Chloride dissolving in water.)

CaCl2  Ca2+ + 2 Cl-

I need an answer that explains the steps
1
Expert's answer
2020-05-27T13:35:45-0400

Cm=m/Mr x V.= 3/111 x 0.5 = 0.054 M.

0.054 x 2 = 0.108 M Cl-.